OpenStudy (anonymous):

When a 58 g tennis ball is served, it accelerates from rest to a constant speed of 36 m/s. The impact with the racket gives the ball a constant acceleration over a distance of 35 cm. What is the magnitude of the net force acting on the ball?

OpenStudy (lalaly):

f=ma we need to find a first v=at= 36m/s x=0.36 m x=1/2 at^2 0.36=1/2(at)t 0.36=1/2 (36) t t=0.02 seconds v=at 36=a*0.02 a=1800 F=ma F=0.058 * 1800 F=104.4 N